Shift report 20010814 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Aug 14 2001 - 09:09:07 EDT

  • Next message: Apache: "Shift report 20010814 08:00-16:00"

    ------------------------------------------------------------
    Date: Tue Aug 14 09:09:07 2001, Shift: 00:00-08:00
    Supervisor: YK Lee, with: M. Murray
    Log book#7 pages 79 - 87
    
    Shift Summary:
    ------------------------------------------------------------
    1.Beam was stable at 100 GeV, 56 bunch of low intensity
    until quench interlock problem at 4 AM.
    2. Run 4168 started from the previous shift did not have
    drift chambers on.
    Run 4169 has ~ 90K events. Run 4172 has ~24 K events.
    3. HV supply L1458A crashed once.
    T1 and T2 have operating voltage of 1200V, and standby
    voltage of 900V.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Aug 14 2001 - 09:09:41 EDT